OpenStudy (anonymous):

(03.02) If g(x) = 2(x - 4), find the value of x if g(x) = 20.

OpenStudy (aum):

g(x) = 2(x - 4) g(x) = 20 20 = 2(x - 4) solve for x.

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

thanks

OpenStudy (aum):

20 = 2(x-4) To solve for x, isolate x. Divide both sides by 2: 10 = x - 4 add 4: 14 = x or x = 14.